Use the following table to answer questions 1 – 6. The wacc is 10% for all projects in this table. Text Effects Year Project5. Projects B and D are mutually exclusive and can be replicated: a. Complete the table below: Project B Project D WACC 10.0%Please help with 5 and 6
Answer #1
Question 5 (a):
Step 1:
Calculation of Equivalent Annual Annuity of Projects B and D
Year Project B Project D
Cash Flows Discount Rate @10% Discounted Cash flows Cash Flows Discount Rate @10% Discounted Cash flows
0 -1000 1 -1000 -1000 1 -1000
1 300 0.909090909 272.7272727 600 0.909090909 545.4545455
2 400 0.826446281 330.5785124 800 0.826446281 661.1570248
3 500 0.751314801 375.6574005
4 600 0.683013455 409.8080732
NPV 388.7713 206.6116
Equivalent Annual Annuity of Project B = NPV / Annuity Discount Factor
   = 388.7713 / 4.16985446
   = 93.23
Equivalent Annual Annuity of Project D = NPV / Annuity Discount Factor
   = 206.6116 /2.73553719
   = 75.53
Step 2:
Calculation of NPV of Projects B and D using Replacement Chain Method
In this case project B should be repeated 3 times and project D for 5 times
So that projects will have equal life
Year Project B Project D
Cash Flows Discount Rate @10% Discounted Cash flows Cash Flows Discount Rate @10% Discounted Cash flows
0 -1000 1 -1000 -1000 1 -1000
1 300 0.909090909 272.7272727 600 0.909090909 545.4545455
2 400 0.826446281 330.5785124 800 0.826446281 661.1570248
3 500 0.751314801 375.6574005 -1000 0.751314801 -751.3148009
4 600 0.683013455 409.8080732 600 0.683013455 409.8080732
5 -1000 0.620921323 -620.9213231 800 0.620921323 496.7370584
6 300 0.56447393 169.342179 -1000 0.56447393 -564.4739301
7 400 0.513158118 205.2632473 600 0.513158118 307.8948709
8 500 0.46650738 233.2536901 800 0.46650738 373.2059042
9 600 0.424097618 254.458571 -1000 0.424097618 -424.0976184
10 -1000 0.385543289 -385.5432894 600 0.385543289 231.3259737
11 300 0.350493899 105.1481698 800 0.350493899 280.3951196
12 400 0.318630818 127.4523271 -1000 0.318630818 -318.6308177
13 500 0.28966438 144.8321899 600 0.28966438 173.7986278
14 600 0.263331254 157.9987526 800 0.263331254 210.6650034
NPV 780.0557731 631.9250345
Step 3:
Project B Project D
WACC 10% 10%
Equivalent Annual Annuity 93.23 75.53
Replacement Chain (RC) NPV 780.06 631.93

Question 5b:

Project B should be selected as Equivalent Annual Annuity and Replacement Chain (RC) NPV for Project B is higher than Project D

Step 1:
Year Project A Project B Project C Project D
Cash Flows Discount Rate @10% Discounted Cash flows Cash Flows Discount Rate @10% Discounted Cash flows Cash Flows Discount Rate @10% Discounted Cash flows Cash Flows Discount Rate @10% Discounted Cash flows
0 -1000 1 -1000 -1000 1 -1000 -1000 1 -1000 -1000 1 -1000
1 1000 0.909090909 909.0909091 300 0.909090909 272.7272727 550 0.909090909 500 600 0.909090909 545.4545455
2 0.826446281 0 400 0.826446281 330.5785124 450 0.826446281 371.9008264 800 0.826446281 661.1570248
3 0.751314801 0 500 0.751314801 375.6574005 350 0.751314801 262.9601803
4 0.683013455 0 600 0.683013455 409.8080732 250 0.683013455 170.7533638
NPV -90.9091 388.7713 305.6144 206.6116
Step 2:
Equivalent Annual Annuity of Project A = NPV / Annuity Discount Factor
   = -90.9091 /1.909091
   = -47.14
Equivalent Annual Annuity of Project B = NPV / Annuity Discount Factor
   = 388.7713 / 4.16985446
   = 93.23
Equivalent Annual Annuity of Project C = NPV / Annuity Discount Factor
   = 305.6144 /4.16985446
   = 73.29
Equivalent Annual Annuity of Project D = NPV / Annuity Discount Factor
   = 206.6116 /2.73553719
   = 75.53
Step 3:
Project B and C has higher NPV hence these should be selected based on NPV rule
Project B and D has higher Equivalent Annual Annuity is higher hence these should be selected based on EAA rule